ameliaxx:

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half . 1) y − 6 = one half (x + 3) 2) y = one half x − 6 3) y + 3 = one half (x − 6) 4) x − 2y = 12

XxAngelGirlxX:

The point-slope form of the equation the line that passes through the point (a,b) and has the slope m is y-b = m(x-a) Therefore, the equation the line that passes through the point (6, -3) and has the slope 1/2 is y - (-3) = 1/2 (x - 6) y + 3 = 1/2 (x-6)

XxAngelGirlxX:

please close this question when done <3